# SplatCo: Structure-View Collaborative Gaussian Splatting for Detail-Preserving Rendering of Large-Scale Unbounded Scenes

### v3

@@ IV-A Rendering Loss
IV-B Structural Similarity Index Measure Loss
@@
IV-C Fine-Grained Gaussian Regularization Loss
@@ IV-D Overall Loss
V Experiments
V-A Datasets
@@
Public Scenes. We perform evaluations on eight scenes, which include the Rubble and Building from the Mill-19 dataset [67], MatrixCity from the city-scale MatrixCity-Aerial dataset [74], and Train and Truck from the Tanks & Temples dataset [75]. Additionally, we select three particularly challenging aerial scenes from areas 1, 4, and 5 of the WHU dataset [76] to assess the performance of various methods.
@@ V-B Evaluation Metrics
Following previous methods [7, 67], we primarily evaluate the quality of the rendered images using three metrics: PSNR, SSIM [73], and the VGG implementation of LPIPS [78].
V-C Implementation Details
@@
Our experiments are conducted on an NVIDIA 3090 GPU using PyTorch, with each scene trained for 30,000 iterations. In the cross-structure collaboration module, we progressively activate the process of acquiring cross-structure features at three granularities—coarse to fine—during training iterations 1, 12,000, and 21,000, respectively. At each level, the pixel resolution of the tri-plane feature maps is doubled compared to the previous level, while the density of the context grid is also increased by a factor of two. To optimize both computational efficiency and rendering performance, we apply the tri-plane spatial attention mechanism for enhancement solely at the first level’s tri-plane feature maps. In the cross-view assisted training strategy, we sample four pairs of images and camera parameters in each training iteration. Additionally, we set the hyperparameter $\tau$ to 1.0 to maximize rendering performance. In our experiments, the MLPs used to update Gaussian attributes, as well as those in the cross-structure collaboration module, both consist of two layers, with the latter incorporating an additional batch normalization operation. All parameters related to Gaussian optimization and the densification process are set according to the original configurations from previous methods [71].
@@
Figure 4: Novel view synthesis results on the Tanks & Temples dataset. Our method, leveraging structure-view assisted learning, achieves superior detail in challenging scenarios. This is evidenced by the sharper reconstruction of distant mountain edges in the train scene and the comprehensive modeling of glass reflection details in the truck scene, highlighting the efficacy of our approach.
V-D Result Analysis
V-D1 Result on Mill-19 and MatrixCity dataset
@@
We rigorously evaluate our method through comprehensive quantitative and qualitative analyses. As shown in Table. I, our approach achieves state-of-the-art performance across three scenarios: MatrixCity, Rubble and Building. In the Building scene, our approach achieves substantial gains over Scaffold-GS, with margins of +0.048 (SSIM), +1.48 (PSNR), and +0.058 (LPIPS), underscoring its superior rendering quality. We attribute this superiority to our collaborative framework’s enhanced adaptability to spatially structured scenes - compared to Scaffold-GS’s hash grid encoding, our architecture models geometric priors more effectively. Notably, Pixel-GS exhibits marked performance degradation in this scenario, suggesting that pixel-optimized point cloud processing struggles with large-scale environments containing complex spatial structures. For a fair comparison, we utilize the official open-source code of MVGS, which is based on the 3DGS framework. The sampled camera-image pair parameter $M$ in MVGS is also set to 4, maintaining consistency with our method. Our method demonstrably outperforms MVGS across the MatrixCity, Rubble, and Building scenes in most metrics. In the Rubble scene only, our method achieves an LPIPS score slightly below 0.012. The superior performance is primarily attributed to the Structure-View Collaboration Optimization approach, which mitigates the detail learning degradation issue in large-scale scenes,effectively preserving the ability to learn both the overall layout and local details. Fig. 3 presents visual comparisons between our method and existing 3DGS baselines. The experimental results show that our approach consistently preserves high-fidelity detail reconstruction while effectively mitigating artifacts common in other methods. Particularly in the Building scene, our method achieves more accurate reconstruction of fine structural details such as solar panel grid patterns, demonstrating clear advantages over competing approaches.
@@
TABLE II: Quantitative comparison on Tanks & Temples dataset. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
@@
[…]
Mip-NerF360 [CVPR22] [79] 0.660 19.52 0.354 0.857 24.91 0.159
iNPG [ToG22] [80] 0.666 20.17 0.386 0.779 23.26 0.274
Plenoxels [CVPR22] [81] 0.663 18.93 0.422 0.774 23.22 0.335
3DGS [ToG23] [7] 0.815 22.07 0.208 0.882 25.46 0.147
GaMeS [arXiv24] [69] 0.816 22.11 0.205 0.877 25.24 0.151
Compact3DGS [CVPR24] [70] 0.792 21.56 0.240 0.871 25.07 0.163
Scaffold-GS [CVPR24] [71] 0.816 22.32 0.213 0.886 25.83 0.141
Pixel-GS [ECCV24] [72] 0.813 22.01 0.188 0.872 25.03 0.131
MVGS [arxiv24] [13] 0.833 22.43 0.155 0.881 25.45 0.116
@@
TABLE III: Quantitative Comparison on SCUT_CA outdoor scenes. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
@@
3DGS [ToG23] [7] 0.887 29.87 0.143 0.849 28.54 0.189 0.882 29.12 0.132
GaMeS [arXiv24] [69] 0.891 29.99 0.140 0.861 28.83 0.182 0.889 29.36 0.120
Compact3DGS [CVPR24] [70] 0.873 29.03 0.164 0.844 27.96 0.211 0.863 27.97 0.160
Scaffold-GS [CVPR24] [71] 0.880 29.49 0.155 0.854 28.44 0.197 0.873 28.67 0.147
Pixel-GS [ECCV24] [72] 0.915 30.81 0.097 0.882 29.34 0.142 0.909 29.77 0.092
@@
V-D2 Result on Tanks & Temples dataset
We present the quantitative results on Tanks and Temples in Table. II. Our method demonstrates superiority over other competitive approaches across most metrics on the ’train’ scene. Specifically, compared to Scaffold-GS, our approach achieves significant improvements of 0.024 in SSIM, 1.44 in PSNR, and 0.055 in LPIPS. For a fair comparison, the camera-image pair sampling parameter $M$ in MVGS is also set to 4. Our method achieves performance gains over MVGS, yielding improvements of $0.007$ and $1.33$ on the SSIM and PSNR metrics, respectively. We also note that our method falls marginally behind MVGS on the LPIPS metric for the ’train’ and ’Truck’ scenes. As visually confirmed in Fig. 4, our method effectively separates the background sky from distant mountains in the rendered imagery. For the ’truck’ scene, our method exhibits superior performance in both PSNR and LPIPS metrics compared to other approaches. While it shows a marginal 0.005 deficit in SSIM relative to Scaffold-GS, our method produces visually superior results. Specifically, it better reconstructs fire hydrant contours and significantly reduces artifacts in the truck’s window regions.
V-D3 Result on SCUT-CA scenes
We further evaluate our method’s performance on our custom SCUT-CA dataset. As shown in Table. III, our approach demonstrates superior performance across all evaluation metrics in both the School History Museum and EI School scenes when compared to other baseline methods. While exhibiting a marginal 0.003 deficit in LPIPS when compared to Pixel-GS, our method shows significant advantages in both SSIM and PSNR metrics. Notably, Scaffold-GS exhibits substantial performance degradation on our custom dataset in comparison to public scene datasets, whereas our method maintains consistently robust performance. This enhanced robustness can be attributed to our structure-view collaborative learning mechanism. The comparative results in Fig. 5 reveal that the original 3DGS fails to reconstruct the zebra crossing details in the scene, while our method not only accurately recovers these fine details but also significantly reduces rendering artifacts. Furthermore, we observe that in the School History Museum scene, both 3DGS and Compact3DGS fail to effectively render vehicle wheels.
TABLE IV: Quantitative Comparison on WHU dataset and plateau region scenes. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
@@ V-D3 Result on SCUT-CA scenes
3DGS [ToG23] [7] 0.887 30.84 0.216 0.901 28.87 0.200 0.799 26.04 0.329 0.723 28.46 0.334 0.745 29.62 0.369
GaMeS [arXiv24] [69] 0.902 31.28 0.207 0.892 28.34 0.203 0.812 26.48 0.311 0.783 29.24 0.308 0.759 29.79 0.363
Compact3DGS [CVPR24] [70] 0.922 32.69 0.217 0.946 33.79 0.147 0.865 28.37 0.273 0.721 28.32 0.355 0.721 29.78 0.400
Scaffold-GS [CVPR24] [71] 0.978 38.90 0.068 0.980 38.86 0.063 0.964 34.70 0.089 0.787 29.66 0.308 0.737 30.06 0.393
Pixel-GS [ECCV24] [72] 0.814 23.51 0.343 0.806 21.07 0.321 0.631 17.61 0.474 0.663 28.20 0.486 0.743 30.42 0.440
@@
V-D4 Result on WHU dataset and Plateau Region scenes
We systematically evaluate our method on both the WHU dataset and plateau scenes, with quantitative and qualitative results presented in Table. IV and Fig. 6, respectively. The WHU dataset is characterized by significant image blurring, while the plateau scenes exhibit distinct planar features. As demonstrated in Table. IV, our method maintains superior performance metrics even under challenging low-quality blurry conditions. Notably, in the plateau scene evaluation, our method achieves optimal results in Scene 2 but exhibits marginally degraded performance in Scene 1. Through rigorous analysis, we attribute the performance gap to non-uniform sampling during data acquisition in Scene 1, where the drone captured images at varying altitudes during its ascent. This inconsistency in viewpoint distribution disrupts the cross-view collaborative learning mechanism that our proposed Cross-View Assisted Training Strategy relies on, ultimately leading to degraded model performance. Furthermore, our experiments reveal that mesh-based methods (e.g., GaMeS) demonstrate particularly strong performance in planar regions, owing to their inherent advantages in geometric representation using mesh structures. Visual comparisons in Fig. 6 show that our method does not have artifacts, and contains more details with much higher accuracy than other methods.
V-E Ablation Study
V-E1 Effectiveness of Cross-Structure Collaboration Module
We conduct an ablation study on our proposed cross-structure collaboration module under three configurations: (1) using only context grid encoding, (2) using only tri-plane feature encoding, and (3) combining both encodings. Additionally, we compare two feature fusion strategies: direct concatenation and attention-based fusion. As shown in Table. V, our experiments demonstrate that (1) fusing both encoding structures significantly improves performance, and (2) employing a lightweight attention mechanism further enhances the results. Qualitative results in Fig. 7 show that our cross-structure collaboration module generates rendered results with sharper line details.
TABLE V: Ablation study on cross-structure collaborated module.
@@ V-E1 Effectiveness of Cross-Structure Collaboration Module
Figure 7: Impact of the cross-structure collaborated module, demonstrated via ablation. Red and green boxes demarcate visual distinctions. By fusing attention-enhanced tri-plane features with context grid features, the module yields demonstrably more refined rendering and improved structural integrity. This stems from its enhanced ability to perceive both overarching point cloud architecture and fine-grained local details. Please zoom in to view detailed results.
V-E2 Effectiveness of Cross-View Assisted Training Strategy
In this section, we conduct ablation studies on the MVRL module, MVAD module, and CVCL module. All modules are implemented using our cross-view assisted training strategy. Furthermore, we specifically optimize the CVCL module through pruning operations. As shown in Table. VI, integrating these key components enables the model to achieve optimal performance, with particularly significant improvements in SSIM metrics. Qualitative results in Fig. 8 demonstrate that incorporating these components significantly enhances rendering quality by improving both detail clarity (e.g., street lamp) and artifact suppression.
TABLE VI: Ablation study on cross-view assisted training strategy.

### v4

@@ 4.1 Rendering Loss
4.2 Structural Similarity Index Measure Loss
@@
4.3 Fine-Grained Gaussian Regularization Loss
@@ 4.4 Overall Loss
5 Experiments
Table 1: Quantitative Comparison on Mill-19 and MatrixCity datasets. The ’-’ symbol represents Mega-NeRF [exp_meganerf] and Switch-NeRF [exp_switchnerf] were not evaluated on MatrixCity. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
MatrixCity Rubble Building
[…]
Mega-NeRF [CVPR22] [exp_meganerf]
0.553 24.06 0.516 0.547 20.93 0.504
Switch-NeRF [ICLR22] [exp_switchnerf]
0.562 24.31 0.496 0.579 21.54 0.474
3DGS [ToG23] [relate_3dgs] 0.737 23.79 0.384 0.746 24.30 0.324 0.676 20.11 0.372
GaMeS [arXiv24] [exp_games] 0.773 24.64 0.367 0.767 24.64 0.311 0.698 20.97 0.329
Compact3DGS [CVPR24] [exp_compact3dgs] 0.795 26.03 0.370 0.719 23.01 0.373 0.645 20.00 0.425
Scaffold-GS [CVPR24] [relate_scaffoldgs] 0.842 27.25 0.290 0.736 23.71 0.346 0.711 21.55 0.328
Pixel-GS [ECCV24] [exp_pixelgs] 0.814 25.60 0.324 0.771 24.67 0.308 0.700 21.00 0.352
MVGS [arxiv24] [mvgs] 0.836 26.92 0.271 0.772 25.82 0.268 0.731 22.46 0.272
CityGS-X [ICCV25] [citygs-x] 0.828 26.40 0.309 0.772 24.96 0.290 0.739 22.20 0.278
Ours 0.855 28.05 0.249 0.794 26.90 0.280 0.759 23.03 0.270
5.1 Datasets
@@
Public Scenes. We perform evaluations on eight scenes, which include the Rubble and Building from the Mill-19 dataset [exp_meganerf], MatrixCity from the city-scale MatrixCity-Aerial dataset [exp_matrixcity], and Train and Truck from the Tanks & Temples dataset [exp_tnt]. Additionally, we select three particularly challenging aerial scenes from areas 1, 4, and 5 of the WHU dataset [exp_whu] to assess the performance of various methods.
@@ 5.2 Evaluation Metrics
Following previous methods [relate_3dgs, exp_meganerf], we primarily evaluate the quality of the rendered images using three metrics: PSNR, SSIM [exp_sim], and the VGG implementation of LPIPS [exp_llism].
5.3 Implementation Details
@@
Our experiments are conducted on an NVIDIA 3090 GPU using PyTorch, with each scene trained for 30,000 iterations. In the CSCM, we progressively activate the process of acquiring HDE features at three granularities—coarse to fine—during training iterations 1, 12,000, and 21,000, respectively. At each level, the pixel resolution of the tri-plane feature maps is doubled compared to the previous level, while the density of the context grid is also increased by a factor of two. To optimize both computational efficiency and rendering performance, we apply the tri-plane spatial attention mechanism for enhancement solely at the first level’s tri-plane feature maps. In our experiments, the MLPs used to update Gaussian attributes, as well as those in the CSCM, both consist of two layers, with the latter incorporating an additional batch normalization operation. All parameters related to Gaussian optimization and the densification process are set according to the original configurations from previous methods [relate_scaffoldgs].
@@
Figure 4: Novel view synthesis results on the Tanks & Temples dataset. Our method, leveraging the SVC module, achieves superior detail in challenging scenarios. This is evidenced by the sharper reconstruction of distant mountain edges in the train scene and the comprehensive modeling of glass reflection details in the truck scene, highlighting the efficacy of our approach.
5.4 Result Analysis
5.4.1 Result on Mill-19 and MatrixCity dataset
@@
We rigorously evaluate our method through comprehensive quantitative and qualitative analyses. As shown in Table. 1, our approach achieves state-of-the-art performance across three scenarios: MatrixCity, Rubble and Building. In the Building scene, our approach achieves substantial gains over Scaffold-GS, with margins of +0.048 (SSIM), +1.48 (PSNR), and +0.058 (LPIPS), underscoring its superior rendering quality. We attribute this superiority to our collaborative framework’s enhanced adaptability to spatially structured scenes - compared to Scaffold-GS’s hash grid encoding, our architecture models geometric priors more effectively. Notably, Pixel-GS exhibits marked performance degradation in this scenario, suggesting that pixel-optimized point cloud processing struggles with large-scale environments containing complex spatial structures. For a fair comparison, we utilize the official open-source code of MVGS, which is based on the 3DGS framework. The sampled camera-image pair parameter $M$ in MVGS is also set to 4, maintaining consistency with our method. Our method demonstrably outperforms MVGS across the MatrixCity, Rubble, and Building scenes in most metrics. In the Rubble scene only, our method achieves an LPIPS score slightly below 0.012. The superior performance is primarily attributed to our SVC module, which mitigates the detail learning degradation issue in large-scale scenes, effectively preserving the ability to learn both the overall layout and local details. For a fair comparison, we also use the open-source code of CityGS-X and train it for the same number of iterations. Our method demonstrates superior performance to CityGS-X across all metrics. Fig. 3 presents visual comparisons between our method and existing 3DGS baselines. The experimental results show that our approach consistently preserves high-fidelity detail reconstruction while effectively mitigating artifacts common in other methods. Particularly in the Building scene, our method achieves more accurate reconstruction of fine structural details such as solar panel grid patterns, demonstrating clear advantages over competing approaches.
@@
Table 2: Quantitative comparison on Tanks & Temples dataset. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
@@
Metrics SSIM PSNR LPIPS SSIM PSNR LPIPS
Mip-NerF360 [relate_mipnerf360] 0.660 19.52 0.354 0.857 24.91 0.159
iNPG [exp_ingp] 0.666 20.17 0.386 0.779 23.26 0.274
Plenoxels [exp_plenoxels] 0.663 18.93 0.422 0.774 23.22 0.335
3DGS [relate_3dgs] 0.815 22.07 0.208 0.882 25.46 0.147
GaMeS [exp_games] 0.816 22.11 0.205 0.877 25.24 0.151
Compact3DGS [exp_compact3dgs] 0.792 21.56 0.240 0.871 25.07 0.163
Scaffold-GS [relate_scaffoldgs] 0.816 22.32 0.213 0.886 25.83 0.141
Pixel-GS [exp_pixelgs] 0.813 22.01 0.188 0.872 25.03 0.131
MVGS [mvgs] 0.833 22.43 0.155 0.881 25.45 0.116
@@
Table 3: Quantitative Comparison on SCUT_CA outdoor scenes. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
@@
3DGS [ToG23] [relate_3dgs] 0.887 29.87 0.143 0.849 28.54 0.189 0.882 29.12 0.132
GaMeS [arXiv24] [exp_games] 0.891 29.99 0.140 0.861 28.83 0.182 0.889 29.36 0.120
Compact3DGS [CVPR24] [exp_compact3dgs] 0.873 29.03 0.164 0.844 27.96 0.211 0.863 27.97 0.160
Scaffold-GS [CVPR24] [relate_scaffoldgs] 0.880 29.49 0.155 0.854 28.44 0.197 0.873 28.67 0.147
Pixel-GS [ECCV24] [exp_pixelgs] 0.915 30.81 0.097 0.882 29.34 0.142 0.909 29.77 0.092
@@
5.4.2 Result on Tanks & Temples dataset
We present the quantitative results on Tanks and Temples in Table. 2. Our method demonstrates superiority over other competitive approaches across most metrics on the ’Train’ scene. Specifically, compared to Scaffold-GS, our approach achieves significant improvements of 0.024 in SSIM, 1.44 in PSNR, and 0.055 in LPIPS. For a fair comparison, the camera-image pair sampling parameter $M$ in MVGS is also set to 4. Our method achieves performance gains over MVGS, yielding improvements of $0.007$ and $1.33$ on the SSIM and PSNR metrics, respectively. We also note that our method falls marginally behind MVGS on the LPIPS metric for the ’Train’ and ’Truck’ scenes. As visually confirmed in Fig. 4, our method effectively separates the background sky from distant mountains in the rendered imagery. For the ’Truck’ scene, our method exhibits superior performance in both PSNR and LPIPS metrics compared to other approaches. While it shows a marginal 0.005 deficit in SSIM relative to Scaffold-GS, our method produces visually superior results. Specifically, it better reconstructs fire hydrant contours and significantly reduces artifacts in the truck’s window regions.
5.4.3 Result on SCUT-CA scenes
We further evaluate our method’s performance on our custom SCUT-CA dataset. As shown in Table. 3, our approach demonstrates superior performance across all evaluation metrics in both the School History Museum and EI School scenes when compared to other baseline methods. While exhibiting a marginal 0.003 deficit in LPIPS when compared to Pixel-GS, our method shows significant advantages in both SSIM and PSNR metrics. Notably, Scaffold-GS exhibits substantial performance degradation on our custom dataset in comparison to public scene datasets, whereas our method maintains consistently robust performance. This enhanced robustness can be attributed to our SVC module. The comparative results in Fig. 5 reveal that the original 3DGS fails to reconstruct the zebra crossing details in the scene, while our method not only accurately recovers these fine details but also significantly reduces rendering artifacts. Furthermore, we observe that in the School History Museum scene, both 3DGS and Compact3DGS fail to effectively render vehicle wheels.
Table 4: Quantitative Comparison on WHU dataset and plateau region scenes. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
@@ 5.4.3 Result on SCUT-CA scenes
3DGS [ToG23] [relate_3dgs] 0.887 30.84 0.216 0.901 28.87 0.200 0.799 26.04 0.329 0.723 28.46 0.334 0.745 29.62 0.369
GaMeS [arXiv24] [exp_games] 0.902 31.28 0.207 0.892 28.34 0.203 0.812 26.48 0.311 0.783 29.24 0.308 0.759 29.79 0.363
Compact3DGS [CVPR24] [exp_compact3dgs] 0.922 32.69 0.217 0.946 33.79 0.147 0.865 28.37 0.273 0.721 28.32 0.355 0.721 29.78 0.400
Scaffold-GS [CVPR24] [relate_scaffoldgs] 0.978 38.90 0.068 0.980 38.86 0.063 0.964 34.70 0.089 0.787 29.66 0.308 0.737 30.06 0.393
Pixel-GS [ECCV24] [exp_pixelgs] 0.814 23.51 0.343 0.806 21.07 0.321 0.631 17.61 0.474 0.663 28.20 0.486 0.743 30.42 0.440
@@
5.4.4 Result on WHU dataset and Plateau Region scenes
We systematically evaluate our method on both the WHU dataset and plateau scenes, with quantitative and qualitative results presented in Table. 4 and Fig. 6, respectively. The WHU dataset is characterized by significant image blurring, while the plateau scenes exhibit distinct planar features. As demonstrated in Table. 4, our method maintains superior performance metrics even under challenging low-quality blurry conditions. Notably, in the plateau scene evaluation, our method achieves optimal results in Scene 2 but exhibits marginally degraded performance in Scene 1. Through rigorous analysis, we attribute the performance gap to non-uniform sampling during data acquisition in Scene 1, where the drone captured images at varying altitudes during its ascent. This inconsistency in viewpoint distribution disrupts the structure-view collaborative learning mechanism that our proposed SVC module relies on, ultimately leading to degraded model performance. Furthermore, our experiments reveal that mesh-based methods (e.g. GaMeS) demonstrate particularly strong performance in planar regions, owing to their inherent advantages in geometric representation using mesh structures. Visual comparisons in Fig. 6 show that our method does not have artifacts, and contains more details with much higher accuracy than other methods.
Table 5: Ablation study on cross-structure collaborated module.
@@ 5.4.4 Result on WHU dataset and Plateau Region scenes
Figure 7: Impact of the CSCM, demonstrated via ablation. Red and green boxes demarcate visual distinctions. By fusing attention-enhanced tri-plane features with context grid features, the module yields demonstrably more refined rendering and improved structural integrity. This stems from its enhanced ability to perceive both overarching point cloud architecture and fine-grained local details. Please zoom in to view detailed results.
5.5 Ablation Study
5.5.1 Effectiveness of Cross-Structure Collaboration Module
We conduct an ablation study on our proposed CSCM under three configurations: (1) using only context grid encoding, (2) using only tri-plane feature encoding, and (3) combining both encodings. Additionally, we compare two feature fusion strategies: direct concatenation and attention-based fusion. As shown in Table. 5, our experiments demonstrate that (1) fusing both encoding structures significantly improves performance, and (2) employing a lightweight attention mechanism further enhances the results. Qualitative results in Fig. 7 show that our CSCM generates rendered results with sharper line details.
Table 6: Ablation study on Cross-View Pruning Mechanism.